OpenStudy (anonymous):

HELP? 194= 1/2x + 2(x-8)

OpenStudy (anonymous):

distribute the 2(x-8): 194=1/2x+2x-16 add the 16 to both sides: 210=1/2x + 2x add the x components: 210= 2.5x (if its (1/2)x but if its 1/(2x ) it will require a different methor divide 2.5 by both sides

OpenStudy (anonymous):

x=210/2.5=84

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

has you solve that one?

OpenStudy (anonymous):

@uzumakhi Yes

OpenStudy (anonymous):

so where is x in numinator or denominator